OpenStudy (anonymous):

In a standard television set, the screen height is 0.75 times the screen width. If a television set measures 33 inches diagonal, what is the screen width? I know that the answer is 26.4 but I don't know how to get to that answer. Please help and thank you! :)

OpenStudy (anonymous):

It has something to do with the Pythagorean theory.

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

Yes i do know that equation but how do i input the numbers?

OpenStudy (anonymous):

let the screen width equal= x so height equal 0.75x we know that (0.75x)^2+x^2= the diagonal d^2 so: 1.5625x^2=33^2 and x^2= 696.96 therefore x= 26.4 which is the width good luck, anything you didn't get let me know!

OpenStudy (anonymous):

ohhhh :D I understand now. Thank you very much!!

OpenStudy (anonymous):

yeh just make sure you understand the initial assumptions about the width and height of the tv, and how they are related.

OpenStudy (anonymous):

got it :)